# Code-as-Symbolic-Planner: Foundation Model-Based Robot Planning via Symbolic Code Generation

### v1

@@ Abstract
Recent works have shown great potentials of Large Language Models (LLMs) in robot task and motion planning (TAMP). Current LLM approaches generate text- or code-based reasoning chains with sub-goals and action plans. However, they do not fully leverage LLMs’ symbolic computing and code generation capabilities. Many robot TAMP tasks involve complex optimization under multiple constraints, where pure textual reasoning is insufficient. While augmenting LLMs with predefined solvers and planners improves performance, it lacks generalization across tasks. Given LLMs’ growing coding proficiency, we enhance their TAMP capabilities by steering them to generate code as symbolic planners for optimization and constraint verification. Unlike prior work that uses code to interface with robot action modules, we steer LLMs to generate code as solvers, planners, and checkers for TAMP tasks requiring symbolic computing, while still leveraging textual reasoning to incorporate common sense. With a multi-round guidance and answer evolution framework, the proposed Code-as-Symbolic-Planner improves success rates by average 24.1% over best baseline methods across seven typical TAMP tasks and three popular LLMs. Code-as-Symbolic-Planner shows strong effectiveness and generalizability across discrete and continuous environments, 2D/3D simulations and real-world settings, as well as single- and multi-robot tasks with diverse requirements. See our project website [^2] https://yongchao98.github.io/Code-Symbol-Planner/ for prompts, videos, and code.
@@ I INTRODUCTION
Pre-trained Large Language Models (LLMs) have shown remarkable performance on many language-related tasks [4], driving a surge of interest in applying them to robot TAMP [5, 6]. A common approach [7, 4, 8, 9, 10] uses LLMs to select actions from pre-defined skill primitives, completing tasks step by step with text or code as intermediates, such as SayCan [11], Inner Monologue [12], Code-as-Policy [13], and ProgGPT [14]. However, these methods rely solely on the LLM’s common-sense reasoning, without leveraging symbolic computation, making them unreliable for complex TAMP tasks involving numeric constraints, optimization, search, and logic. Another line of work integrates LLMs with traditional TAMP planners, such as Text2Motion [5] and AutoTAMP [15], where LLMs translate natural language instructions into formal representations for downstream TAMP solvers. While these methods incorporate symbolic reasoning, their generalizability is limited, as each task requires carefully designed planners and customized integration frameworks.
To address the limitations of existing LLM-based TAMP approaches, we propose a new paradigm: querying LLMs to directly generate symbolic code that serves as both the robot’s TAMP planner and checker. We refer to this approach as Code-as-Symbolic-Planner. Unlike Code-as-Policy [13], which uses LLMs to generate code as intermediates for connecting with action modules — relying only on commonsense reasoning, Code-as-Symbolic-Planner generates symbolic code that explicitly performs efficient search and reasoning over valid plans, incorporating symbolic computation into the planning process. Since LLMs can flexibly generate diverse codes for different tasks, this approach naturally integrates symbolic computation into TAMP while preserving broad generalizability.
@@ II RELATED WORK
Task and Motion Planning Robotic planning spans both high-level discrete task planning [16] and low-level continuous motion planning [17], collectively known as task and motion planning [1]. Existing approaches typically follow one of three strategies: (1) ensuring motion feasibility before sequencing actions [18, 19], (2) generating action sequences first, then resolving motion constraints [20, 21], or (3) interleaving task and motion reasoning [22, 23]. However, these methods are inherently limited to goals expressible using predefined predicates and executable through a fixed set of symbolic operators [24, 25].
@@
LLMs for Code Generation Modern LLMs are extensively trained on diverse code datasets [31, 32]. Recent work has focused on optimizing agent frameworks, improving training strategies, and simplifying tasks to better utilize LLMs as capable software developers [33, 34, 35]. Another line of research explores using LLMs to generate code for solving mathematical and logical problems [36, 37]. In this work, we investigate the potential and challenges of using LLM-generated code to solve TAMP tasks.
@@ III PROBLEM DESCRIPTION
The goal is to convert natural language instructions — including spatial, logical, and temporal constraints — into task plans (sequences of pre-defined actions) or motion plans (sets of timed waypoints, e.g., ($x_{i},y_{i},t_{i}$)). For motion planning, the environment is represented as named obstacles described by polygons with locations and orientations. Drones, racecars, and robot arms must stay within maximum velocity limits, and total operation time must not exceed the task’s time limit. Full motion trajectories are assumed to be linear interpolations between waypoints, with complex trajectories achievable through denser waypoint sequences.
@@ V EXPERIMENTAL DESIGN
We compare the LLM-based planning methods on four discrete task planning tasks and three continuous motion planning tasks including both single- and multi-robot settings, as shown in Figure 3 and 4, respectively. In task planning scenarios, such as BoxNet, Blocksworld, BoxLift, and Gridworld, LLMs are required to generate sequences of high-level actions. In motion planning scenarios, such as Path-Racecars, Shape Formation, and Path-Drones, LLMs are required to generate sequences of motion waypoints. In addition to the 2D scenarios for our experiments, we also perform experiments in a 3D environment simulated using Pybullet [39] and real hardware settings with two robot arms, as shown in Figure 4 and 7. These tested tasks originate from previous works [40, 28, 41, 15].
@@
Each task comprises 140 testing samples with varied complexities and constraints. We evaluate the success of each testing trial by checking three criteria: (1) syntax correctness, (2) task completion, and (3) satisfaction of required constraints, such as collision avoidance, time limits, and logical consistency. For each method, the LLM is prompted with a task description without any few-shot examples. If the generated code takes more than 50 seconds to execute, we terminate the trial and record it as a failure. The specific descriptions of the testing tasks are as follows:
@@
BoxNet In Figure 3(a), the environment consists of cell regions, robot arms, colored boxes, and corresponding colored goal locations (circles). The objective is to move each box to its matching goal location in the fewest time steps. Each robot arm is restricted to the cell it occupies and can perform two actions: (1) move a box to a neighboring cell or (2) place a box in a goal location within the same cell.
@@
Blocksworld In Figure 3(b), the goal is to stack blocks in a specified order. The robot can take four actions: (1) pick up a block, (2) unstack a block from another block, (3) put down a block, and (4) stack a block on another block.
@@
BoxLift In Figure 3(c), robots are tasked with lifting boxes in the fewest time steps. Each robot has a different lifting capacity, and each box has a different weight. Multiple robots can collaborate to lift the same box in a single time step. A box is successfully lifted when the combined lifting capacity of the assigned robots exceeds the box’s weight.
@@
Gridworld Figure 3(d) consists of obstacles (black) and goals (red). The robot needs to visit all goals, and any attempt to move into obstacles or move out of the grid will result in failure. The robot has five possible actions: (1) move up, (2) move down, (3) move left, (4) move right, (5) visit goal.
@@
Path-Racecars Figure 4(a) queries LLMs to plan the racecar trajectory waypoints to move all the racecars from one side to another under varied environments and obstacles.
@@
Shape Formation Figure 4(b) queries LLMs to plan the picking order and placing positions of randomly located colored boxes to form required shapes like circles and triangles. The locations and shapes of bowls are unchangeable so that the boxes should not occupy bowls’ areas.
@@
Path-Drones Figure 4(c) requires LLMs to plan the drone trajectory waypoints to pass through the hole one by one while keeping safe distances to the wall and other drones.
@@ VI RESULTS
Notably, Code-as-Symbolic-Planner shows the best scalability on more complex tasks compared to the other three baselines. Ideally, the symbolic code it generates should handle these tasks effectively. However, in some highly complex cases, the generated code exceeds the 50-second execution time limit, contributing to its performance drop in the most challenging scenarios.
@@ VII CONCLUSION
Limitations and Future Directions First, while our results show that current LLMs can generate correct code to solve TAMP tasks under multi-round guidance, the generated code is often not optimal, especially for high-complexity tasks where it may exceed time limits. In robotics, many existing planners are highly efficient but lack generalizability. Combining Code-as-Symbolic-Planner with these efficient planners is a promising direction for future work.

### v2

@@ Abstract
Recent works have shown great potential of Large Language Models (LLMs) in robot task and motion planning (TAMP). Current LLM approaches generate text- or code-based reasoning chains with sub-goals and action plans. However, they do not fully leverage LLMs’ symbolic computing and code generation capabilities. Many robot TAMP tasks involve complex optimization under multiple constraints, where pure textual reasoning is insufficient. While augmenting LLMs with predefined solvers and planners improves performance, it lacks generalization across tasks. Given LLMs’ growing coding proficiency, we enhance their TAMP capabilities by steering them to generate code as symbolic planners for optimization and constraint verification. Unlike prior work that uses code to interface with robot action modules or pre-designed planners, we steer LLMs to generate code as solvers, planners, and checkers for TAMP tasks requiring symbolic computing, while still leveraging textual reasoning to incorporate common sense. With a multi-round guidance and answer evolution framework, the proposed Code-as-Symbolic-Planner improves success rates by average 24.1% over best baseline methods across seven typical TAMP tasks and three popular LLMs. Code-as-Symbolic-Planner shows strong effectiveness and generalizability across discrete and continuous environments, 2D/3D simulations and real-world settings, as well as single- and multi-robot tasks with diverse requirements. See our project website [^2] https://yongchao98.github.io/Code-Symbol-Planner/ for prompts, videos, and code.
@@ I INTRODUCTION
Pre-trained Large Language Models (LLMs) have shown remarkable performance on many language-related tasks [4], driving a surge of interest in applying them to robot TAMP [5, 6]. A common approach [4, 7, 8, 9] uses LLMs to select actions from pre-defined skill primitives, completing tasks step by step with text or code as intermediates, such as SayCan [10], Inner Monologue [11], Code-as-Policies [12], and ProgPrompt [13]. However, these methods rely solely on the LLM’s common-sense reasoning, without leveraging symbolic computation, making them unreliable for complex TAMP tasks involving numeric constraints, optimization, search, and logic. Another line of work integrates LLMs with traditional TAMP planners, such as Text2Motion [5] and AutoTAMP [14], where LLMs translate natural language instructions into formal representations for downstream TAMP solvers. While these methods incorporate symbolic reasoning, their generalizability is limited, as each task requires carefully designed planners and customized integration frameworks.
To address the limitations of existing LLM-based TAMP approaches, we propose a new paradigm: querying LLMs to directly generate code that serves as both the robot’s TAMP planner and checker. We refer to this approach as Code-as-Symbolic-Planner. Unlike Code-as-Policies [12], which uses LLMs to generate code as intermediates for connecting with action modules—relying primarily on commonsense reasoning without efficient symbolic computation, Code-as-Symbolic-Planner generates code that explicitly performs efficient search and reasoning over valid plans, incorporating symbolic computation into the planning process. Since LLMs can flexibly generate diverse codes for different tasks, this approach naturally integrates symbolic computation into TAMP while preserving broad generalizability.
@@ II RELATED WORK
Task and Motion Planning Robotic planning spans both high-level discrete task planning [15] and low-level continuous motion planning [16], collectively known as task and motion planning [1]. Existing approaches typically follow one of three strategies: (1) ensuring motion feasibility before sequencing actions [17], (2) generating action sequences first, then resolving motion constraints [18, 19], or (3) interleaving task and motion reasoning [20]. However, these methods are inherently limited to goals expressible using predefined predicates and executable through a fixed set of symbolic operators [21, 22].
@@
LLMs for Code Generation Modern LLMs are extensively trained on diverse code datasets [28, 29]. Recent work has focused on optimizing agent frameworks, improving training strategies, and simplifying tasks to better utilize LLMs as capable software developers [30, 31]. Another line of research explores using LLMs to generate code for solving mathematical and logical problems [32, 33]. In this work, we investigate the potential and challenges of using LLM-generated code to solve TAMP tasks.
@@ III PROBLEM DESCRIPTION
The goal is to convert natural language instructions—including spatial, logical, and temporal constraints—into task plans (sequences of pre-defined actions) or motion plans (sets of timed waypoints, e.g., ($x_{i},y_{i},t_{i}$)). For motion planning, the environment is represented as named obstacles described by polygons with locations and orientations. Drones, racecars, and robot arms must stay within maximum velocity limits, and total operation time must not exceed the task’s time limit. Full motion trajectories are assumed to be linear interpolations between waypoints, with complex trajectories achievable through denser waypoint sequences. The planned output of LLMs should be the sequence of waypoints and timepoints for each robot or drone.
@@ V EXPERIMENTAL DESIGN
We compare the LLM-based planning methods on four discrete task planning tasks and three continuous motion planning tasks including both single- and multi-robot settings, as shown in Figure 3 and 4, respectively. In task planning scenarios, such as BoxNet, Blocksworld, BoxLift, and Gridworld, LLMs are required to generate sequences of high-level actions. In motion planning scenarios, such as Path-Racecars, Shape Formation, and Path-Drones, LLMs are required to generate sequences of motion waypoints. In addition to the 2D scenarios for our experiments, we also perform experiments in a 3D environment simulated using Pybullet [35] and real hardware settings with two robot arms, as shown in Figure 4 and 7. These tested tasks originate from previous works [36, 25, 37, 14, 38, 39].
@@
Each task includes 140 test samples with diverse complexities and constraints. To evaluate the method’s ability to handle highly complex tasks, the environment is intentionally made more challenging than in previous studies. We evaluate the success of each testing trial by checking three criteria: (1) syntax correctness, (2) task completion, and (3) satisfaction of required constraints, such as collision avoidance, time limits, and logical consistency. For each method, the LLM is prompted with a task description without any few-shot examples. If the generated code takes more than 50 seconds to execute, we terminate the trial and record it as a failure. The specific descriptions of the testing tasks are as follows:
@@
BoxNet [25] In Figure 3(a), the environment consists of cell regions, robot arms, colored boxes, and corresponding colored goal locations. The objective is to move each box to its matching goal location in the fewest time steps. Each robot arm is restricted to the cell it occupies and can perform two actions: (1) move a box to a neighboring cell or (2) place a box in a goal location within the same cell.
@@
Blocksworld [40] In Figure 3(b), the goal is to stack blocks in a specified order. The robot can take four actions: (1) pick up a block, (2) unstack a block from another block, (3) put down a block, and (4) stack a block on another block.
@@
BoxLift [25] In Figure 3(c), robots are tasked with lifting boxes in the fewest time steps. Each robot has a different lifting capacity, and each box has a different weight. Multiple robots can collaborate to lift the same box in a single time step. A box is successfully lifted when the combined lifting capacity of the assigned robots exceeds the box’s weight.
@@
Gridworld [38] Figure 3(d) consists of obstacles (black) and goals (red). The robot needs to visit all goals. Attempts to move into obstacles or move out of the grid will result in failure. The robot has five possible actions: (1) move up, (2) move down, (3) move left, (4) move right, (5) visit goal.
@@
Path-Racecars [14] Figure 4(a) queries LLMs to plan the racecar trajectory waypoints to move all the racecars from one side to another under varied environments and obstacles.
@@
Shape Formation [39] Figure 4(b) queries LLMs to plan the picking order and placing positions of randomly located colored boxes to form required shapes like circles and triangles. The locations and shapes of bowls are unchangeable so that the boxes should not occupy bowls’ areas.
@@
Path-Drones [14] Figure 4(c) requires LLMs to plan the drone trajectory waypoints to pass through the hole one by one while keeping safe distances to the wall and other drones.
@@ VI RESULTS
Notably, Code-as-Symbolic-Planner shows the best scalability on more complex tasks compared to the other three baselines. Ideally, the symbolic code it generates should handle these tasks effectively.
Failure Reasons of Code-as-Symbolic-Planner However, in some highly complex cases, the generated code exceeds the 50-second execution time limit, contributing to its performance drop in the most challenging scenarios. This also demonstrates one typical failure case of Code-as-Symbolic-Planner that sometimes the generated code does not utilize the most efficient algorithms or the task is too complex so that purely algorithm-based methods are not sufficient. Another major bottleneck we identified is the LLM’s inability to translate real-world application problems into abstract optimization or logistical tasks solvable through code, which is also challenging to human experts.
@@ VII CONCLUSION
Limitations and Future Directions First, while our results show that current LLMs can generate correct code to solve TAMP tasks under multi-round guidance, the generated code is often not optimal, especially for high-complexity tasks where it may exceed time limits. Generating planners with coding also risks adding errors. In robotics, many existing planners are highly efficient but lack generalizability. Combining Code-as-Symbolic-Planner with these efficient planners is a promising direction for future work.